OpenStudy (anonymous):

There are two points on the graph of y=x^3 where the tangent lines are parallel to y=x. Find these points

OpenStudy (anonymous):

dy1/dx1 = dy2/dx2 ,

OpenStudy (anonymous):

differentiate both equations implicitly with respect to x and set them equal to eachother

OpenStudy (anonymous):

ok. I'm going to see if I can work it out

OpenStudy (anonymous):

Will the answer be d/dx=3x^4 ?

OpenStudy (anonymous):

3x^2, you minus for differentiation

OpenStudy (anonymous):

so you would have dy1/dx1=3x^2 and dy2/dx2=1

OpenStudy (anonymous):

yeah that's what I meant..lol

OpenStudy (anonymous):

i thought as much :)

OpenStudy (anonymous):

you can quickly see that you will indeed have two answers because of the squared x term

OpenStudy (anonymous):

You got dy2/dx2 by the x from the parallel tangent of the line right?

OpenStudy (anonymous):

Are the two points 3 and 6? I got 6 for the second point by finding the second derivative

OpenStudy (anonymous):

the 2, is just a subscript, sorry for the confusion. Parrallel lines have equal slopes, and the derivative is a slope so by solving fot them being equal to eachother you will find where they are parallel.

OpenStudy (anonymous):

\[3x^2=1 \implies x^2 = 1/3 \implies x= +- \sqrt(1/3)\]

OpenStudy (anonymous):

Oh ok. Just looked it over myself and definitely got it. Thanks so much for your help!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

you have to plug those x values back into your original equation to get the points!

OpenStudy (anonymous):

oh. I thought we were done..lol. So I plugged it back into the y=x^3 equation and got .19 and -.19

OpenStudy (anonymous):

it askes for the points, so you need (x,y) we have x so we need y. we get y by plugin in the points. It should be -1/27, 1/27

OpenStudy (anonymous):

but for x didn't we get the +- square root of 1/3 and not just 1/3

OpenStudy (anonymous):

oops your right! sorry http://www.wolframalpha.com/input/?i=%28sqrt%281%2F3%29%29%5E3

OpenStudy (anonymous):

damnit i got an imaginary number http://www.wolframalpha.com/input/?i=%28sqrt%28-1%2F3%29%29%5E3

OpenStudy (anonymous):

lol. Is that a good or bad thing? Clearly, I need to brush up on my math

OpenStudy (anonymous):

I mean it's still a number